Operations Report – January 7th

SOL: 6
Name of person filing report: Mason Kuhn
Non-nominal systems: Some old-generation EVA Radio Headsets
Notes on non-nominal systems: After the EVA started today, 2 crew members were experiencing comms issues. Before departing on the rovers, those 2 crew members returned to the Hab to troubleshoot the issues. It was determined the radios had no issues, and all that was required was a different headset. Later in the night, all old-generation headsets will be tested to determine functionality (please see ‘Summary of RAM operations’ for additional information).
Spirit rover used: No
Hours: 206.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Opportunity rover used: No
Hours: 111.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 218.7
Beginning charge: 100%
Ending charge: 73%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 253.5
Beginning charge: 100%
Ending charge: 69%
Currently charging: Yes
General notes on rovers: Curiosity and Perseverance were used on today’s EVA, and the performance of each was nominal.
Summary of Hab operations: Four crew members took showers today, resulting in higher but still reasonable/sustainable daily water usage.
WATER USE: 26.1 gallons
Water (static tank): 272.3
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: No
Summary of internet: Nominal
Summary of suits and radios: Suits 6, 8, 9, and 11 were used on the EVA. I do not currently have an update on suit 7 due to circumstances of yesterday evening, but the Commander and I will investigate suit 7 airflow issues tonight. All radios are nominal, but not all older-generation headsets are nominal. We received new headsets from an interplanetary cargo ship, tested all of them, and all are nominal.
Summary of GreenHab operations: Watered and monitored all GreenHab crops. Watered experimental crops. Planted more microgreens.
WATER USE: 7 gallons
Heater: On, door propped open & monitored for temperature control
Supplemental light: 4 hours
Harvest: 10g dill, 1g basil, 1g chives, 20g microgreens
Summary of ScienceDome operations:
Dual split: On (from ~ 8:00 pm until 8:00 am)
Summary of RAM operations: Retrieved the new EVA Radio Headsets from the RAM and brought them to the lower deck. All have been tested, verified to be nominal, and labeled. Later tonight, I will organize and label the old-generation headsets into nominal or faulty categories & place them back in the RAM.
Summary of any observatory issues: N/A
Summary of health and safety issues: See HSO Covid Test Report.
Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site